Determine gm in the circuit  if IDSS=10mA and VGS(off)= -8V .

1.41mS

2.5mS

undefined

1.875mS

+ 20 ha 301 A 076
Answer #1

Ip ss = loma Tos Tossal - Wapis vp = Vas (Off) = -8 V Tos = lom(1-6) alom (po) Ips = 5.625mA gm = Eps loss (11 56 25m) lom) [